# Sparse-to-Sparse Training of Diffusion Models

### v1

@@ 1 Introduction
Diffusion models (DMs) are a class of deep generative models that exhibit extraordinary performance to produce diverse and high-quality data. DMs currently dominate the generative field in computer vision, having been applied to a wide range of tasks such as (un)conditional image generation [Ho et al., 2020b, Rombach et al., 2021, Nichol and Dhariwal, 2021, Dhariwal and Nichol, 2021, Nichol et al., 2022, Blattmann et al., 2022, Das et al., 2023], image super-resolution [Saharia et al., 2021, Chung et al., 2022], and image inpainting [Nichol et al., 2022, Chung et al., 2022, Saharia et al., 2022], among others. DMs have also shown incredible potential in other domains, including speech generation [Liu et al., 2023a], text generation [Li et al., 2022, Gong et al., 2023], and time-series prediction and imputation [Rasul et al., 2021, Tashiro et al., 2021].
@@
Despite these advantages, DMs are notorious for their slow training, demanding significant computational resources and resulting in a considerable carbon footprint [Strubell et al., 2020]. Due to the extensive number of diffusion timesteps required to produce a single sample (e.g., Rombach et al. [2021] mentioned up to 500 steps), DMs also suffer from slow sampling speed [Song et al., 2021]. Even though progress has been made in improving inference speed, DMs are still considerably slower than other generative approaches such as GANs and VAEs [Rombach et al., 2021]. This inefficiency impacts not only end users, but also the research community, by hindering further developments due to the lengthy process of model training and evaluation.
@@
Reducing the computational costs and memory requirements of DMs is a critical challenge for the broad implementation and adoption of these models, and an active field of research. Much of the existent literature has addressed this challenge through improvements to the inference stage [Song et al., 2021, Nichol and Dhariwal, 2021, Fang et al., 2023, Shang et al., 2023, Li et al., 2023, Salimans and Ho, 2022, Meng et al., 2023]. Efforts have also been made in the direction of training efficiency, exploring different architectures and training strategies [Wang et al., 2023, Ding et al., 2023, Rombach et al., 2021, Phung et al., 2022], but training DMs is still an extensive and costly process.
@@
In the last few years, sparse-to-sparse training has emerged as a promising approach to significantly reduce the computational cost of deep learning models, by training sparse networks from scratch [Mocanu et al., 2018, Bellec et al., 2018, Dettmers and Zettlemoyer, 2019, Evci et al., 2020, Zhang et al., 2024b]. Interestingly, sparse neural networks have been shown to match, or even outperform, their Dense counterparts in classification tasks [Mocanu et al., 2018, Liu et al., 2021a], generative modeling using GANs [Liu et al., 2023b], and Reinforcement Learning [Sokar et al., 2022], all while requiring less memory and reducing the number of floating-point operations (FLOPs). We should note that, currently, most sparse neural networks require roughly the same amount of time to train as their dense counterparts, since today’s hardware is optimized for dense matrix operations. However, growing interest in sparse models is reshaping the landscape; see Appendix A for a discussion in this regard.
@@
We propose to lower the computational cost of DMs by incorporating, for the first time, the paradigm of sparse-to-sparse training for unconditional generation. As such, we introduce three different methods, Static-DM (static strategy), RigL-DM, and MagRan-DM (both dynamic strategies), that can be easily integrated with existing DMs. Since our goal is to study the effect of these techniques on the performance of DMs, we experiment using two state-of-the-art DMs in two domains: Latent Diffusion [Rombach et al., 2021] for image generation (continuous, pixel-level data) and ChiroDiff [Das et al., 2023] for sketch generation (discrete, spatiotemporal sequence data).
In sum, we make the following contributions:
@@
  • We introduce sparse-to-sparse training to DMs, with both static and dynamic strategies. We consider various sparsity levels, two state-of-the-art models (Latent Diffusion and ChiroDiff), and six datasets in total.
@@ 2.1 Diffusion Models
DMs [Sohl-Dickstein et al., 2015, Ho et al., 2020a, Song et al., 2020] are probabilistic models designed to learn a data distribution $q(x)$ through two processes: a forward noising process and a reverse denoising process. The forward process is defined as a Markov Chain of length $T$ in which Gaussian noise is added at each timestep $t$, producing a sequence of increasingly noisier samples:
@@
The reverse denoising process $q(x_{t-1}|x_{t})$, attempts to recover the original data, but it is intractable as it depends on the entire data distribution $q(x)$. As such, we need to parameterize a neural network $p_{\theta}$ to approximate it. This network $p_{\theta}$ can be optimized by training with the simplified objective, $\mathcal{L}=\mathbb{E}_{t\sim[1,T],x_{0},\epsilon\sim\mathcal{N}(0,1)}||\epsilon-\epsilon_{\theta}(x_{t},t)||^{2}$, where $x_{T}$ is a noisy version of input $x$ at the final timestep $T$, and $\epsilon_{\theta}$ the prediction of the neural network $p_{\theta}$.
@@ 2.2 Efficiency in Diffusion Models
Increasing the efficiency of DMs has been primarily addressed through accelerating the sampling process, by reducing the number of diffusion steps through faster sampling [Song et al., 2021, Karras et al., 2022] and model distillation [Salimans and Ho, 2022, Meng et al., 2023, Yin et al., 2024]. As for training acceleration, some works have proposed shifting the diffusion process to the latent space [Rombach et al., 2021, Vahdat et al., 2021]. Interestingly, Phung et al. [2022] used discrete wavelet transforms to decompose images into sub-bands, employing these sub-bands to perform the diffusion more efficiently.
Previous studies have also presented refinements to the training process of DMs. For example, Wang et al. [2023] introduced a plug-and-play training strategy that utilizes patches instead of the full images, to improve training speed. Hang et al. [2024] proposed treating DMs as a multitask learning problem and introduced a weighting strategy to balance the different timesteps, achieving a significant improvement in training convergence speed.
From the perspective of network compression, prior works have explored techniques such as structural pruning [Fang et al., 2023], post-training quantization [Shang et al., 2023, Li et al., 2023], knowledge distillation [Yang et al., 2023], and the lottery ticket hypothesis [Frankle and Carbin, 2019, Jiang et al., 2023]. Very recently, Wang et al. [2024] proposed the incorporation of sparse masks into pre-trained DMs before fine-tuning, and achieved a 50% reduction in multiply-accumulate operations (MACs) with only a slight average decrease of image quality (as measured by the FID score). Although these techniques work in increasing efficiency, they still require pre-training of full DMs. Our work proposes training sparse DMs from scratch, which has the potential to both accelerate training and inference, and reduce the memory footprint.
@@ 2.3 Sparse-to-Sparse Training
In SST methods, the connectivity pattern between neurons is set at initialization, and remains fixed during training. This concept was first introduced by Mocanu et al. [2016], who proposed a non-uniform scale-free topology for Restricted Boltzmann machines, with the sparse models achieving better results than their Dense counterparts. Later, Liu et al. [2022] investigated the efficacy of random pruning at initialization, and found that, using appropriate layer-wise sparsity ratios, a randomly pruned subnetwork of WideResNet-50 can outperform a dense WideResNet-50 on ImageNet. Many other criteria have been proposed to set layer-wise sparsity ratios before training, by trying to identify important connections using information such as connection sensitivity, as in SNIP [Lee et al., 2019], gradient flow [Wang et al., 2020], as in GraSP. Very recently, two new initialization criteria have been proposed that utilize concepts from network science theory: Bipartite Scale-Free and Bipartite Small-World [Zhang et al., 2024a, b].
@@
In DST methods, the network is initialized with a connectivity pattern and dynamically explores different connections throughout training [Mocanu et al., 2018, Bellec et al., 2018]. This was first proposed by Mocanu et al. [2018] through Sparse Evolutionary Training (SET), an algorithm that adjusts the connections using a prune-and-grow scheme every $N$ training steps. In SET, weights are dropped based on their magnitude (ensuring an equal amount of positive and negative weights) and regrown randomly. RigL [Evci et al., 2020] proposes an alternative method that prunes the weights based on the absolute magnitude, and regrows them based on the gradients by calculating the dense gradients only at the update step. Although further pruning methods have been proposed [Lee et al., 2019, Yuan et al., 2021], a study by Nowak et al. [2023] found only minor differences between the tested criteria. The contrast was higher in lower density patterns, with magnitude pruning giving the best performance. Other growing criteria have been proposed based on randomness [Mostafa and Wang, 2019] and momentum [Dettmers and Zettlemoyer, 2019].
Recently, Zhang et al. [2024b] proposed Epitopological Sparse Meta-deep Learning (ESML), a brain-inspired, gradient-free method, which aims to shift the focus from the weights to the network topology, and uses concepts from network theory. By leveraging ESML, the authors train a sparse network that using just 1% of the connections, is able to surpass dense networks, as well as other DST methods, in several image classification tasks.
DST has also been applied to the field of generative modelling: Liu et al. [2023b] proposed STU-GAN, comprised of a generator with high sparsity and a denser discriminator. STU-GAN was able to outperform a dense BigGAN on CIFAR-10 with a 80% sparse generator and 70% sparse discriminator.
@@ 3 Methodology
Our study aims to understand the effect of sparse-to-sparse training techniques on DMs. We focus on unstructured sparsity due to its ability to maintain high performance even at very high levels of sparsity [Evci et al., 2020]. Thus, our experiments cannot rely on current hardware to accelerate sparse computations; for example, NVIDIA A100 and Ampere cards only support 2:4 structured sparsity, which requires to enforce a fixed sparsity level of 50%. In the following sections, we present three methods of introducing sparsity in DMs: one SST technique, Static-DM, and two DST techniques, MagRan-DM and RigL-DM.
@@ 3.1 Static Sparse Training: Static-DM
1:  Input: Dataset $\mathcal{D}$, Network $f_{\theta}$, Number of Epochs $N$, Diffusion steps $T_{d}$, Sparsity ratio $S$
2:  $\theta\leftarrow$ sparse initialization using $S$
3:  for $i=1$ to $N$ do
@@
Following the findings of Liu et al. [2022], we randomly prune the connections at initialization using the Erdõs–Rényi (ER) [Mocanu et al., 2018] strategy to allocate the non-zero weights to non-convolutional layers. With this strategy, larger layers get assigned higher sparsity than smaller layers. The sparsity of each layer scales with $s^{l}\propto 1-\frac{n^{l}+n^{l-1}}{n^{l}\cdot n^{l-1}}$, where $n^{l}$ and $n^{l-1}$ represent the number of neurons in layer $l$ and $l-1$ respectively.
For convolutional layers, we use a modification of ER, ERK [Evci et al., 2020], which takes into account the size of the kernels, $s^{l}\propto 1-\frac{n^{l}+n^{l-1}+w^{l}+h^{l}}{n^{l}\cdot n^{l-1}\cdot w^{l}\cdot h^{l}}$, where $n^{l}$ and $n^{l-1}$ represent the number of neurons in layer $l$ and $l-1$ respectively, and $w^{l}$ and $h^{l}$ the width and height of the corresponding convolutional kernel.
@@ 3.2 Dynamic Sparse Training: MagRan-DM and RigL-DM
The key aspect of DST algorithms lies with the process of pruning and regrowing weights. We opted to test the two most common regrowth methods, random growth and gradient growth, combined with the magnitude pruning criteria. Magnitude pruning is a simple criteria, that has been shown to perform well in high sparsity regimes for supervised classification, as well as in other generative models [Nowak et al., 2023, Liu et al., 2023b]
RigL, proposed by Evci et al. [2020], combines gradient growth and magnitude pruning, thus the name of our model RigL-DM. The combination of random growth and magnitude pruning closely resembles the SET algorithm [Mocanu et al., 2018], and has been studied before for other types of models [Nowak et al., 2023], although it has never been named. For simplicity, we refer to this method as MagRan-DM.
@@
1:  Input: Dataset $\mathcal{D}$, Network $f_{\theta}$, Number of Epochs $N$, Diffusion steps $T_{d}$, Sparsity ratio $S$, exploration frequency $\Delta T_{e}$, Pruning rate $p$, Sparse method method
2:  $\theta\leftarrow$ sparse initialization using $S$
3:  for $i=1$ to $N$ do
@@
9:        $\theta_{i_{p}}$ = TopMag(|$\absolutevalue{\theta_{i}}$ |, $1-p$) // Magnitude pruning
@@
11:           $\theta_{i_{g}}$ = TopGrad(|$\absolutevalue{\triangledown_{\theta}\mathcal{L}_{\text{DIF}}}$ |, $p$) // Gradient growth
@@ 3.3.1 Models and Benchmarks
We test Static-DM, MagRan-DM, and RigL-DM against the Dense baseline, on two different DMs, Latent Diffusion [Rombach et al., 2021] and ChiroDiff [Das et al., 2023], on the task of unconditional image generation. Although image generation is the most common application and main direction of current research in DMs, we seek to offer a more extensive look, and examined DMs for different modalities, with different backbone architectures. More detailed information about the model architectures and choice of datasets can be found in Appendix B.
@@
Latent Diffusion is a DM that creates high-quality images while reducing computational requirements by training in a compressed lower-dimensional latent space. Although we focus on unconditional generation tasks, Latent Diffusion also allows for conditional generation, by using a general-purpose mechanism based on cross attention [Vaswani et al., 2017]. Latent Diffusion first employs pre-trained autoencoders to obtain a latent representation of the input, and then performs the diffusion process on these representations, using a U-Net [Ronneberger et al., 2015]. Performing the denoising process in the latent space allows to the model to focus on relevant semantic-wise information about the data. We sparsify only the U-Net model, as shown in Figure 1a, and utilize off-the shelf autoencoders provided by Rombach et al. [2021], keeping them dense. We evaluate on the LSUN-Bedrooms [Yu et al., 2015], CelebA-HQ [Karras et al., 2018] and Imagenette [Howard, 2019] datasets.
@@
ChiroDiff is a DM specifically designed to model continuous-time chirographic data, such as sketches or handwriting, in the form of a sequence of strokes containing both spatial and temporal information. ChiroDiff can handle sequences of variable length and, as a non-autoregressive model, is able to capture holistic concepts, leading to higher quality samples. This model employs a Bidirectional GRU encoder as backbone architecture. The encoder is fed the spatial coordinates, their point-wise velocities, as well as the entire sequence as context, which provides full context of the sequence during the generation process. Sparsity is applied to the entire network, as shown in Figure 1b. We evaluate it on KanjiVG, QuickDraw [Ha and Eck, 2018], and VMNIST [Das et al., 2022]. Following the original paper, we use a preprocessed version of KanjiVG. [^1] https://github.com/hardmaru/sketch-rnn-datasets/tree/master/kanji For QuickDraw we use the following categories: crab, cat, and mosquito; and all results are averaged.

### v2

@@ 1 Introduction
Diffusion models (DMs) are a class of deep generative models that exhibit extraordinary performance to produce diverse and high-quality data. DMs currently dominate the generative field in computer vision, having been applied to a wide range of tasks such as (un)conditional image generation (Ho et al., 2020b; Rombach et al., 2021; Nichol and Dhariwal, 2021; Dhariwal and Nichol, 2021; Nichol et al., 2022; Blattmann et al., 2022; Das et al., 2023), image super-resolution (Saharia et al., 2021; Chung et al., 2022), and image inpainting (Nichol et al., 2022; Chung et al., 2022; Saharia et al., 2022), among others. DMs have also shown incredible potential in other domains, including speech generation (Liu et al., 2023a), text generation (Li et al., 2022; Gong et al., 2023), and time-series prediction and imputation (Rasul et al., 2021; Tashiro et al., 2021).
@@
Despite these advantages, DMs are notorious for their slow training, demanding significant computational resources and resulting in a considerable carbon footprint (Strubell et al., 2020). Due to the extensive number of diffusion timesteps required to produce a single sample (e.g., Rombach et al. (2021) mentioned up to 500 steps), DMs also suffer from slow sampling speed (Song et al., 2021). Even though progress has been made in improving inference speed, DMs are still considerably slower than other generative approaches such as GANs and VAEs (Rombach et al., 2021). This inefficiency impacts not only end users, but also the research community, by hindering further developments due to the lengthy process of model training and evaluation.
@@
Reducing the computational costs and memory requirements of DMs is a critical challenge for the broad implementation and adoption of these models, and an active field of research. Much of the existent literature has addressed this challenge through improvements to the inference stage (Song et al., 2021; Nichol and Dhariwal, 2021; Fang et al., 2023; Shang et al., 2023; Li et al., 2023; Salimans and Ho, 2022; Meng et al., 2023). Efforts have also been made in the direction of training efficiency, exploring different architectures and training strategies (Wang et al., 2023; Ding et al., 2023; Rombach et al., 2021; Phung et al., 2022), but training DMs is still an extensive and costly process.
@@
In the last few years, sparse-to-sparse training has emerged as a promising approach to significantly reduce the computational cost of deep learning models, by training sparse networks from scratch (Mocanu et al., 2018; Bellec et al., 2018; Dettmers and Zettlemoyer, 2019; Evci et al., 2020; Zhang et al., 2024b). Interestingly, sparse neural networks have been shown to match, or even outperform, their Dense counterparts in classification tasks (Mocanu et al., 2018; Liu et al., 2021a), generative modeling using GANs (Liu et al., 2023b), and Reinforcement Learning (Sokar et al., 2022), all while requiring less memory and reducing the number of floating-point operations (FLOPs). We should note that, currently, most sparse neural networks require roughly the same amount of time to train as their dense counterparts, since today’s hardware is optimized for dense matrix operations. However, growing interest in sparse models is reshaping the landscape; see Appendix A for a discussion in this regard.
@@
We propose to lower the computational cost of DMs by incorporating, for the first time, the paradigm of sparse-to-sparse training for unconditional generation. As such, we introduce three different methods, Static-DM (static strategy), RigL-DM, and MagRan-DM (both dynamic strategies), that can be easily integrated with existing DMs. Since our goal is to study the effect of these techniques on the performance of DMs, we experiment using two state-of-the-art DMs in two domains: Latent Diffusion (Rombach et al., 2021) for image generation (continuous, pixel-level data) and ChiroDiff (Das et al., 2023) for sketch generation (discrete, spatiotemporal sequence data). In sum, we make the following contributions:
@@
  • We introduce sparse-to-sparse training to unconditional DMs, with both static and dynamic strategies. We consider various sparsity levels, two state-of-the-art models (Latent Diffusion and ChiroDiff), and six datasets in total. We also perform some experiments using conditional DMs.
@@ 2.1 Diffusion Models
DMs (Sohl-Dickstein et al., 2015; Ho et al., 2020a; Song et al., 2020) are probabilistic models designed to learn a data distribution $q(x)$ through two processes: a forward noising process and a reverse denoising process. The forward process is defined as a Markov Chain of length $T$ in which Gaussian noise is added at each timestep $t$, producing a sequence of increasingly noisier samples:
@@
[…]
@@ 2.2 Efficiency in Diffusion Models
Increasing the efficiency of DMs has been primarily addressed through accelerating the sampling process, by reducing the number of diffusion steps through faster sampling (Song et al., 2021; Karras et al., 2022) and model distillation (Salimans and Ho, 2022; Meng et al., 2023; Yin et al., 2024). As for training acceleration, some works have proposed shifting the diffusion process to the latent space (Rombach et al., 2021; Vahdat et al., 2021). Interestingly, Phung et al. (2022) used discrete wavelet transforms to decompose images into sub-bands, employing these sub-bands to perform the diffusion more efficiently.
Previous studies have also presented refinements to the training process of DMs. For example, Wang et al. (2023) introduced a plug-and-play training strategy that utilizes patches instead of the full images, to improve training speed. Hang et al. (2024) proposed treating DMs as a multitask learning problem and introduced a weighting strategy to balance the different timesteps, achieving a significant improvement in training convergence speed.
From the perspective of network compression, prior works have explored techniques such as structural pruning (Fang et al., 2023), post-training quantization (Shang et al., 2023; Li et al., 2023), knowledge distillation (Yang et al., 2023), and the lottery ticket hypothesis (Frankle and Carbin, 2019; Jiang et al., 2023). Very recently, Wang et al. (2024) proposed the incorporation of sparse masks into pre-trained DMs before fine-tuning, and achieved a 50% reduction in multiply-accumulate operations (MACs) with only a slight average decrease of image quality (as measured by the FID score). Although these techniques work in increasing efficiency, they still require pre-training of full DMs. Our work proposes training sparse DMs from scratch, which has the potential to both accelerate training and inference, and reduce the memory footprint.
@@ 2.3 Sparse-to-Sparse Training
In SST methods, the connectivity pattern between neurons is set at initialization, and remains fixed during training. This concept was first introduced by Mocanu et al. (2016), who proposed a non-uniform scale-free topology for Restricted Boltzmann machines, with the sparse models achieving better results than their Dense counterparts. Later, Liu et al. (2022) investigated the efficacy of random pruning at initialization, and found that, using appropriate layer-wise sparsity ratios, a randomly pruned subnetwork of WideResNet-50 can outperform a dense WideResNet-50 on ImageNet. Many other criteria have been proposed to set layer-wise sparsity ratios before training, by trying to identify important connections using information such as connection sensitivity, as in SNIP (Lee et al., 2019), gradient flow (Wang et al., 2020), as in GraSP. Very recently, two new initialization criteria have been proposed that utilize concepts from network science theory: Bipartite Scale-Free and Bipartite Small-World (Zhang et al., 2024a; b).
@@
In DST methods, the network is initialized with a connectivity pattern and dynamically explores different connections throughout training (Mocanu et al., 2018; Bellec et al., 2018). This was first proposed by Mocanu et al. (2018) through Sparse Evolutionary Training (SET), an algorithm that adjusts the connections using a prune-and-grow scheme every $N$ training steps. In SET, weights are dropped based on their magnitude (ensuring an equal amount of positive and negative weights) and regrown randomly. RigL (Evci et al., 2020) proposes an alternative method that prunes the weights based on the absolute magnitude, and regrows them based on the gradients by calculating the dense gradients only at the update step. Although further pruning methods have been proposed (Lee et al., 2019; Yuan et al., 2021), a study by Nowak et al. (2023) found only minor differences between the tested criteria. The contrast was higher in lower density patterns, with magnitude pruning giving the best performance. Other growing criteria have been proposed based on randomness (Mostafa and Wang, 2019) and momentum (Dettmers and Zettlemoyer, 2019).
Recently, Zhang et al. (2024b) proposed Epitopological Sparse Meta-deep Learning (ESML), a brain-inspired, gradient-free method, which aims to shift the focus from the weights to the network topology, and uses concepts from network theory. By leveraging ESML, the authors train a sparse network that using just 1% of the connections, is able to surpass dense networks, as well as other DST methods, in several image classification tasks.
DST has also been applied to the field of generative modelling: Liu et al. (2023b) proposed STU-GAN, comprised of a generator with high sparsity and a denser discriminator. STU-GAN was able to outperform a dense BigGAN on CIFAR-10 with a 80% sparse generator and 70% sparse discriminator.
@@ 3 Methodology
Our study aims to understand the effect of sparse-to-sparse training techniques on DMs. We focus on unstructured sparsity due to its ability to maintain high performance even at very high levels of sparsity (Evci et al., 2020). Thus, our experiments cannot rely on current hardware to accelerate sparse computations; for example, NVIDIA A100 and Ampere cards only support 2:4 structured sparsity, which requires to enforce a fixed sparsity level of 50%. In the following sections, we present three methods of introducing sparsity in DMs: one SST technique, Static-DM, and two DST techniques, MagRan-DM and RigL-DM.
@@ 3.1 Static Sparse Training: Static-DM
1:Input: Dataset $\mathcal{D}$, Network $f_{\theta}$, Number of Epochs NN, Diffusion steps $T_{d}$, Sparsity ratio $S$
[…]
3:for $i=1$ to NN do
@@
Following the findings of Liu et al. (2022), we randomly prune the connections at initialization using the Erdõs–Rényi (ER) (Mocanu et al., 2018) strategy to allocate the non-zero weights to non-convolutional layers. With this strategy, larger layers get assigned higher sparsity than smaller layers. The sparsity of each layer scales with $s^{l}\propto 1-\frac{n^{l}+n^{l-1}}{n^{l}\cdot n^{l-1}}$, where $n^{l}$ and $n^{l-1}$ represent the number of neurons in layer $l$ and l1l-1 respectively.
For convolutional layers, we use a modification of ER, ERK (Evci et al., 2020), which takes into account the size of the kernels, $s^{l}\propto 1-\frac{n^{l}+n^{l-1}+w^{l}+h^{l}}{n^{l}\cdot n^{l-1}\cdot w^{l}\cdot h^{l}}$, where $n^{l}$ and $n^{l-1}$ represent the number of neurons in layer $l$ and l1l-1 respectively, and $w^{l}$ and $h^{l}$ the width and height of the corresponding convolutional kernel.
@@ 3.2 Dynamic Sparse Training: MagRan-DM and RigL-DM
The key aspect of DST algorithms lies with the process of pruning and regrowing weights. We opted to test the two most common regrowth methods, random growth and gradient growth, combined with the magnitude pruning criteria. Magnitude pruning is a simple criteria, that has been shown to perform well in high sparsity regimes for supervised classification, as well as in other generative models (Nowak et al., 2023; Liu et al., 2023b)
RigL, proposed by Evci et al. (2020), combines gradient growth and magnitude pruning, thus the name of our model RigL-DM. The combination of random growth and magnitude pruning closely resembles the SET algorithm (Mocanu et al., 2018), and has been studied before for other types of models (Nowak et al., 2023), although it has never been named. For simplicity, we refer to this method as MagRan-DM.
@@
1:Input: Dataset $\mathcal{D}$, Network $f_{\theta}$, Number of Epochs NN, Diffusion steps $T_{d}$, Sparsity ratio $S$, exploration frequency $\Delta T_{e}$, Pruning rate $p$, Sparse method method
[…]
3:for $i=1$ to NN do
@@
9:   $\theta_{i_{p}}$ = TopMag($|\absolutevalue{\theta_{i}}$, 1p1-p) // Magnitude pruning
@@
11:    $\theta_{i_{g}}$ = TopGrad($|\absolutevalue{\triangledown_{\theta}\mathcal{L}_{\text{DIF}}}$, $p$) // Gradient growth
@@ 3.3.1 Models and Benchmarks
We test Static-DM, MagRan-DM, and RigL-DM against the Dense baseline, on two different DMs, Latent Diffusion (Rombach et al., 2021) and ChiroDiff (Das et al., 2023), given their popularity among the research literature, on the task of unconditional image generation. Although image generation is the most common application and main direction of current research in DMs, we seek to offer a more extensive look, and examined DMs for different modalities, with different backbone architectures. More detailed information about the model architectures and choice of datasets can be found in Appendix B.
@@
Latent Diffusion is a DM that creates high-quality images while reducing computational requirements by training in a compressed lower-dimensional latent space. Although we focus on unconditional generation tasks, Latent Diffusion also allows for conditional generation, by using a general-purpose mechanism based on cross attention (Vaswani et al., 2017). Latent Diffusion first employs pre-trained autoencoders to obtain a latent representation of the input, and then performs the diffusion process on these representations, using a U-Net (Ronneberger et al., 2015). Performing the denoising process in the latent space allows to the model to focus on relevant semantic-wise information about the data. We sparsify only the U-Net model, as shown in Figure 1a, and utilize off-the shelf autoencoders provided by Rombach et al. (2021), keeping them dense. We evaluate on the LSUN-Bedrooms (Yu et al., 2015), CelebA-HQ (Karras et al., 2018) and Imagenette (Howard, 2019) datasets.
@@
ChiroDiff is a DM specifically designed to model continuous-time chirographic data, such as sketches or handwriting, in the form of a sequence of strokes containing both spatial and temporal information. ChiroDiff can handle sequences of variable length and, as a non-autoregressive model, is able to capture holistic concepts, leading to higher quality samples. This model employs a Bidirectional GRU encoder as backbone architecture. The encoder is fed the spatial coordinates, their point-wise velocities, as well as the entire sequence as context, which provides full context of the sequence during the generation process. Sparsity is applied to the entire network, as shown in Figure 1b. We evaluate it on KanjiVG, QuickDraw (Ha and Eck, 2018), and VMNIST (Das et al., 2022). Following the original paper, we use a preprocessed version of KanjiVG. [^1] https://github.com/hardmaru/sketch-rnn-datasets/tree/master/kanji For QuickDraw we use the following categories: crab, cat, and mosquito; and all results are averaged.